Hello
So
I have a test on polynomial stuff, factoring, completing the square, factoring trinomials, how to change from one form of an equation (general, vertex, factored form) is there any easy way to remember all this stuff because it seems for each equation you're doing somthing different and I'm so lost... Is there an easy way to remember it all?

Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

I have a test on polynomial stuff, factoring, completing the square, factoring trinomials, how to change from one form of an equation (general, vertex, factored form) is there any easy way to remember all this stuff because it seems for each equation you're doing somthing different and I'm so lost... Is there an easy way to remember it all?
----------------
The only reasonable approach is to do some of the problems.
Look at the solved problems on this site, but don't just look at them.
If you work some, you'll remember what you did, and how to do them.

Answer by ewatrrr(24785)

 
Hi,
As to remembering..it might help to remember WHAT a particular form easily tell us:
Hope this helps.
For ex:
y = x^2 - 6x +8 this tells us the y-intercept when x = 0 is (0,8)
factored: y = (x-4)(x-2) esily gives us the zeros of the functions (when y = 0)
0 = (x-4)(x-2) says x= 4 and x =2 are the x-intercepts
vertex form: y = (x-3)^2 -9 + 8 OR y = 1·(x-3)^2 -1
says Vertex is (3,-1)& as 1>0, parabola opens Upward
We know the graph looks like this: